OpenStudy (anonymous):

Evaluate the integral: (x^(2) + x + 1) / x * dx from e to 1

OpenStudy (anonymous):

\(\Large \int_e ^1 \frac{x^2+x+1}{x}dx = \int_e ^1 x+1+\frac{1}{x}dx \) can you take it from here?

OpenStudy (anonymous):

im not good at integrals :/. I'm confused, so I would really like you to explain how to do it and show the work.

OpenStudy (anonymous):

Like for example, do you take the anti-derivative? do you set U = to something and then take the derivative of that and get rid of DX? I'm kind of confused on the processor

OpenStudy (anonymous):

on the process*

OpenStudy (anonymous):

all i did was simplify the integrand so you'd be able to use the integration shortcuts....

OpenStudy (anonymous):

yeah I know, but I mean in general, how would you solve an integral problem?

OpenStudy (anonymous):

I don't exactly know the steps to doing one is what I'm saying

OpenStudy (anonymous):

ok... let's break up the integral.... \(\Large \int_e ^1 x+1+\frac{1}{x}dx= \int_e ^1xdx + \int_e ^11dx + \int_e ^1\frac{1}{x}dx\) can you find that first integral: \(\Large \int_e ^1xdx= \) ????

OpenStudy (anonymous):

The anti-derivative of X is x^(2) / 2.